# Exploring Spatial Representation to Enhance LLM Reasoning in Aerial Vision-Language Navigation

### v2

@@ IV Experiment
LingUNet [misra2018mapping] 383.8 0.6 6.9 0.2
Seq2Seq [anderson2018vision] 146.0 4.8 19.8 1.6
CMA [CMA] 121.0 3.0 23.2 0.6
AerialVLN [liu_2023_AerialVLN] 90.2 7.2 15.7 2.4
Navid* [navid] 105.1 6.8 15.5 1.1
Openfly* [openfly] 127.2 8.1 21.8 1.6
Ours 96.3 12.6 31.6 2.7
TABLE II: Comparison on the validation unseen of the AerialVLN-S. * denotes we train these methods by running their source codes.
Validation Unseen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Random 149.7 0.0 0.0 0.0
Action Sampling 237.6 0.2 1.1 0.1
LingUNet [misra2018mapping] 368.4 0.4 3.6 0.9
Seq2Seq [anderson2018vision] 218.9 2.3 11.7 0.7
CMA [CMA] 172.1 3.2 16.0 1.1
AerialVLN [liu_2023_AerialVLN] 127.9 5.1 10.5 1.4
Navid* [navid] 106.9 6.1 12.3 0.7
Openfly* [openfly] 113.8 7.6 18.2 1.5
Ours 119.5 10.8 23.0 1.9
TABLE III: Comparison on OpenFly dataset unseen split.
OpenFly Unseen
model NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
Random 301m 0.1 0.1 0
Seq2Seq[anderson2018vision] 291m 0.1 3.2 0.1
CMA[CMA] 268m 0.7 6.6 0.1
AerialVLN[liu_2023_AerialVLN] 227m 2.3 19.8 1.3
Navid[navid] 200m 2.6 16.9 2.2
OpenFly[openfly] 198m 5.1 27.3 2.0
Ours 205m 8.0 31.0 2.4
IV-A Dataset and Implementation Details
IV-A1 Dataset
We conduct experiments using two datasets to assess aerial VLN task performance, i.e., AerialVLN-S benchmark from [liu_2023_AerialVLN] and the recently introduced OpenFly dataset [openfly]. The AerialVLN-S dataset simulates real-world urban environments with over 870 object categories across diverse scenarios such as downtown cities, factories, parks, and villages. It contains 8,446 expert-curated flying paths recorded by AOPA-certified UAV pilots, ensuring trajectory realism and navigational fidelity. For evaluation, we utilize 12 scenes from its validation set and 5 scenes from the validation unseen set, as the official test set remains inaccessible. To enable broader comparative analysis, we introduce the OpenFly unseen dataset, which provides 1,200 test trajectories uniformly sampled from more diverse scenarios. We compare with several state-of-the-art VLN models on both benchmarks, using OpenFly’s test unseen trajectories for generalization assessment. This dual evaluation strategy rigorously tests the performances of these works, establishing comprehensive baselines for aerial VLN research.
IV-A2 Evaluation metrics
To validate the effectiveness of our method, we utilize a comprehensive set of evaluation metrics following [liu_2023_AerialVLN]. We focus on several key aspects, i.e., Navigation Error (NE) quantifying the distance between the UAV’s stopping point and the actual destination, Success Rate (SR) which measures the proportion of navigations that successfully reach the destination within a 20-meter threshold, Oracle Success Rate (OSR), an idealized measure considering any point on the trajectory that comes within 20 meters of the destination as a success, and Success rate weighted by Normalised Dynamic Time Warping (SDTW), which takes into account both the navigation success rate and the similarity between the real paths.
IV-A3 Implementation details
Our framework is implemented in both a simulator and a real environment. In the simulation environments of AirSim and Unreal Engine 4 (UE4), we evaluated the methods using a PC equipped with an Intel i9 12th-generation CPU and Nvidia RTX 4090 GPUs. For the real world environment, we test on a Q250 airframe, carrying an Intel RealSense D435i depth camera, and an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as onboard computer. All methods are run on an external PC (same as the simulation setup) communicating with the onboard computer via a UDP network. For LLM reasoning we use the online API of GPT-4V and GPT-4o, where the default parameters are employed.
IV-B Experimental Results
IV-B1 Quantitative results in simulator
Table I and  Table II compare our method with several other works on the AerialVLN-S dataset. Random and Action Sampling respectively mean sampling actions randomly and according to the action distribution of the training set. LingUNet [misra2018mapping] assumes an agent can see the destination from the beginning, which is transformed into a step-wise paradigm by [liu_2023_AerialVLN]. Seq2Seq [anderson2018vision] is a learning-based recurrent policy, directly predicting the next action. CMA is a baseline model in [liu_2023_AerialVLN], considering attention between multi-modalities. LAG is the main method in [liu_2023_AerialVLN], designing a look-ahead guidance to adjust paths. Navid [navid] is a VLM-based method, achieving great success in the indoor VLN task. OpenFly [openfly] proposes a keyframe-aware aerial VLN model, taking key historical observations into account. In comparison, our method significantly outperforms others in most metrics, demonstrating the superiority of our method. Notably, our method is training-free, while others are trained using corresponding data.  Table III shows the generalization results of different methods on the OpenFly Unseen dataset, where our method also achieves better performance than other works.
We found that instead of processing the entire long instruction and visual observations at each time step, the formatted decomposed sub-goal is more conducive to the fine-grained alignment of each action and landmark. Moreover, UAV has a better understanding of its own spatial position and other semantic constraints under the guidance of our STMR, which has obvious advantages in long instruction environments and leads to more robust execution. Aerial VLN is a challenging task that has not been fully studied, and the overall performance needs to be improved. We believe that our work can bring some insights into this field.
IV-B2 Quantitative results in real environment
@@
To test whether our solution works well in the real world, we collect 10 outdoor scenes including street scenes and forests with ground-truth lengths ranging from 50m to 500m. Then, we evaluate our method using a real UAV and call cloud-hosted LLMs to navigate in these challenging environments. Not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT [chen2024mapgpt] and NavGPT [zhou2024navgpt], are employed and compared in this experiment. Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint, we manually define candidate waypoints for them in this experiment. From demonstration snapshot Fig. 6, we can see that our method can align visual and textual landmarks as well as understand commands. Finally, the UAV reaches the destination successfully.  Table IV also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods. However, we can see that LLM-based methods including ours need more computation time. This is mainly because of the latency of the LLM API. We believe that with the development of LLMs and hardware computing power, the latency will gradually decrease.
Figure 4: Demonstration of the STMR in spatial reasoning.
@@
TABLE IV: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
model SR/$\uparrow$ OSR/$\uparrow$ Time (s)
AerialVLN [liu_2023_AerialVLN] 0 10 82
Navid [navid] 10 10 75
Openfly [openfly] 20 30 63
MapGPT [chen2024mapgpt] 20 20 176
NavGPT [zhou2024navgpt] 10 20 150
@@
Figure 6: Qualitative results in real environment. Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence.
IV-B3 Case analysis
Fig. 4 demonstrates the effectiveness of our STMR in helping spatial reasoning, where the left part shows an intermediate STMR during the navigation process, and the right part shows simplified prompt, questions, and reasoning answers. From Fig. 4 we can see that an LLM can understand and describe the surrounding environment through the STMR. It can even make accurate judgments about objects or areas at a distance. Such an ability is extremely helpful for aerial VLN in vast outdoor environments.
We also summarize two most common failures caused by incorrect planning or execution. The first typical failure is the misunderstanding of ambiguous instructions. For example, as shown in the left part of Fig. 5, sometimes there are a lot of continuous commands without any landmark, such as “turn left, then move right, then go straight”, which have no objects for an LLM to refer to. This case will directly lead to the repeated execution of an action. The second typical error is caused by the inaccuracy of the visual perception. Although modern perception models show strong capability, they still need to be improved in identifying objects from different views. As a result, the key landmark may not be mapped to the STMR, causing the proper action not to be completed as illustrated in the right part of Fig. 5. This observation aligns with our real-world experiment.
TABLE V: Ablation study on different spatial representations.
@@ IV-B3 Case analysis
NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
@@
Ours 87.1 19.4 26.2 5.7
IV-C Ablation Studies
@@
We carry out comprehensive ablation studies to assess the core designs of the proposed method. All ablation experiments are conducted on the randomly selected 100 samples from the AerialVLN-S dataset. Corresponding results are depicted in Table V,  Table VI,  Table VII, and Table VIII.
IV-C1 Different types of spatial information
@@
To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt. In Topo format (row 1 of Table V), we maintain a linguistically formed map that captures the topological relationships between different nodes. Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts. For example, “Place 1 is connected with Places 2, 4, 0, 3”. In Metric format (row 2 of Table V), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt. Each vision observation records the direction and distance of the landmark, for example, “a white building in the left front 10 meters away”. It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still has poorer navigation performance, since for some landmarks such as roads and rivers, UAVs still cannot obtain region-level spatial information based on a certain point distance description. Our STMR combining the semantic, topological, and metric information significantly enhances the UAV’s exploration capability and improves the accuracy in SR and OSR by over 10%.
IV-C2 Direct visual input vs STMR
@@
Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they are capable of direct reasoning with visual and textual prompts. However, it is challenging for them to make correct VLN decisions from sequential images. This may be because of the inability of VLMs in spatial-aware reasoning. As shown in Table VI, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs.
IV-C3 Grid size of STMR
[…]
IV-C4 State update of sub-goal
@@
To ensure that each sub-goal can be executed despite environmental changes, in each iteration, the LLM does not change the original path plan, but updates the status of the current sub-goal including todo, in process, and completed. We compare this strategy with updating the entire multi-step path planning (row 1 of Table VIII) according to the observation in each iteration. This strategy does not notably enhance the Oracle Success Rate (OSR), showing that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively. Nonetheless, it positively impacts the decision-making process, resulting in an increase in the success rate (SR) from 9.0% to 10.8%.
TABLE VI: Ablation study on the different visual prompts for VLMs.
VLM Input RGB STMR NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
@@
$\checkmark$ 119.5 10.8 23.0 2.6
TABLE VII: Ablation study on the grid size $r$ of STMR.
Size NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
$10m$ 304.9 2.0 9.5 0.6
$5m$ 119.5 10.8 23.0 1.9
TABLE VIII: Ablation study on the state update strategies.
Module NE/m $\downarrow$ SR/$\uparrow$ OSR/$\uparrow$ SDTW/$\uparrow$
@@
w state 119.5 10.8 23.0 1.9
V CONCLUSIONS AND FUTURE WORK
This paper addresses the challenging aerial VLN task by introducing an LLM-based training-free framework. To enhance the spatial reasoning ability of LLMs, we propose Semantic-Topo-Metric Representation (STMR). STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information. Taking the proposed STMR as a part of the LLM prompts, we significantly improve the UAV’s navigation capabilities. Our framework achieves state-of-the-art results on the AerialVLN-S and Openfly dataset, demonstrating its effectiveness. In the future, we will explore flexible framework that will interact with humans to eliminate ambiguities in the instructions.

### v3

@@ Introduction
LingUNet (Misra et al. 2018) 383.8 0.6 6.9 0.2
Seq2Seq (Anderson et al. 2018) 146.0 4.8 19.8 1.6
CMA (Hong et al. 2022) 121.0 3.0 23.2 0.6
AerialVLN (Liu et al. 2023c) 90.2 7.2 15.7 2.4
Navid (Zhang et al. 2024a) 105.1 6.8 15.5 1.1
OpenUAV (Wang et al. 2024) 102.8 6.3 17.6 2.0
Ours (training-free) 96.3 12.6 31.6 2.7
Table 2: Comparison on the validation seen set of AerialVLN-S.
Experimental Results
Baseline Models
  • $\bullet$
    Rule-based Methods. Rule-based methods include Random and Action Sampling. In the Random method, the agent randomly chooses actions at each step and continues until either the ‘stop’ action is chosen or the maximum number of steps is reached. The Action Sampling method, on the other hand, first analyzes the statistical properties of the dataset and then samples actions based on the action distribution.
  • $\bullet$
    Learning-based Methods. Learning-based methods include mainstream aerial navigation approaches like LingUNet, Seq2Seq and CMA (Liu et al. 2023c), as well as the state-of-the-art method Navid (Zhang et al. 2024a) and OpenUAV (Wang et al. 2024). Different from other baselines, Navid is a method for the indoor VLN task, so we retrain it on the original AerialVLN dataset.
  • $\bullet$
    LLM-based Methods. For LLM-based methods, we compare two works for the indoor VLN task, i.e., MapGPT (Chen et al. 2024b) and NavGPT (Zhou, Hong, and Wu 2024). To ensure fairness, all these methods are evaluated using the GPT-4o with the same setting.
Quantitative Results in Simulator.
We first analyze performance on the simple navigation tasks. The scenes mainly tests the agent’s ability to locate landmarks and predict actions correctly. Even a short instruction like “fly along the road and stop near the intersection” requires the UAV to accurately understand the scene, recognize the destination, and stop in a proper location. As shown in Table 1, our method outperforms other works by a large margin in terms of SR (+26.8%) and OSR (+22.3%). In addition, our method exhibits a significantly lower navigation error (NE) compared to the other methods, which is reduced by 45.6 meters on average compared with the second-place method, Navid. These indicate that the UAV has a better understanding of its own spatial position and semantic constraints under the guidance of our STMR. In Table 2, the proposed method consistently achieves superior performance on the AerialVLN-S dataset. As the task complexity increases significantly with much longer trajectories and instructions, the success rates of all methods experience a notable decline. The results in Table 2 indicate that Aerial VLN remains a challenging task having not been fully studied, and there is still room for improvement in overall performance. We believe that our work can bring valuable insights into this field.
Quantitative Results in Real Environment.
@@
To test the performance of our solution in the real world, we collected 10 outdoor scenes, including street scenes and forests, with ground-truth lengths ranging from 50m to 200m. Then, We evaluate our method using a real UAV, leveraging cloud-hosted LLMs to navigate through these challenging environments. In this experiment, we employ and compare not only pretrained VLN models but also other LLM-based VLN methods, i.e., MapGPT (Chen et al. 2024b) and NavGPT (Zhou, Hong, and Wu 2024). Notably, MapGPT and NavGPT rely on predefined topological maps to predict the next navigation waypoint. In this experiment, we manually define candidate waypoints for both methods.
As shown in the demonstration snapshot in Fig. 6, our method effectively aligns visual and textual landmarks, understands commands, and successfully guides the UAV to its destination.  Table 3 also shows that the proposed method achieves a much better performance than both pretrained VLN methods and training-free LLM-based methods. However, we can see that LLM-based methods, including ours, need more computation time. This is mainly because of the latency of the LLM API. We believe that with the development of LLMs and hardware computing power, the latency will gradually decrease.
@@
model SR/% $\uparrow$ OSR/% $\uparrow$ Time (s)
AerialVLN (Liu et al. 2023c) 0 10 82
Navid (Zhang et al. 2024a) 10 10 75
MapGPT (Chen et al. 2024b) 20 20 176
NavGPT (Zhou, Hong, and Wu 2024) 10 20 150
@@
Table 3: Quantitative results in real environment. The last column shows the average runtime of aerial VLN procedures.
Figure 6: Visualization of successful and failed examples in the real environments. Short-range movements of the same action are merged into a single long-range movement, as shown in the image sequence. The second row illustrates a failed case, the UAV mistakenly executed the “go straight” action because the visual perceptor failed to recognize the road in the fifth image.
Case Analysis.
Fig. 4 demonstrates the effectiveness of our STMR in helping spatial reasoning, where the left part shows an intermediate STMR during the navigation process, and the right part shows simplified prompt, questions, and reasoning answers. As shown in Fig. 4, an LLM can understand and describe the surrounding environment through the STMR. It can even make accurate judgments about objects or areas at a distance. This ability is particularly valuable for aerial VLN in vast outdoor environments.
Fig. 5 illustrates two of the most common failure cases caused by incorrect planning or execution. The first typical failure arises from the misunderstanding of ambiguous instructions. there are many continuous commands without any landmarks, such as turn left, then move right, then go straight, which lack landmarks for the LLM to reference. This ambiguity often leads to the repeated execution of an action. The second typical error is caused by the inaccuracy of visual perception. Although modern perception models show strong capability, they still require improvement in identifying objects from different views. As a result, the key landmark may not be mapped to the STMR, causing the proper action not to be completed as illustrated in the right part of Fig. 5 and the second row of Fig. 6, showcasing the challenges posed by complex environments.
@@
NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
@@
Ours 88.7 15.0 28.0 3.6
Table 4: Ablation study on different spatial representations.
Ablation Studies
@@
We conduct comprehensive ablation studies to assess the core components of the proposed method. We randomly sample 100 samples from the unseen validation dataset of AerialVLN-S and perform all the ablation experiments. Corresponding results are depicted in Table 4,  Table 5,  Fig. 7, and Table 6.
Different Types of Spatial Information.
@@
To demonstrate the spatial representation ability of STMR, we further compare it with other prompting formats as an LLM’s spatial information prompt. In Topo format (row 1 of Table 4), we maintain a linguistically formed map that captures the topological relationships between different nodes. Each node records textual descriptions of visual observations, and the connectivity between nodes is described using textual prompts. For example, Place 1 is connected with Places 2, 4, 0, 3. In the Metric format (row 2 of Table 4), we arrange the visual observations from 8 different directions in a clockwise order relative to the agent’s current orientation and concatenate them into a single prompt. Each visual observation records the direction and distance of the landmark, for example, a white building in the left front 10 meters away. It is shown that row 2 improves SR by 1.2% (absolute) over row 1 but still demonstrates poorer navigation performance. This is because, for certain landmarks like roads and rivers, UAVs still cannot obtain region-level spatial information based solely on point-distance descriptions. In contrast, our STMR combining the semantic, topological, and metric information, significantly enhances the UAV’s exploration capability and improves OSR accuracy by over 10%.
Direct Visual Input vs STMR.
@@
Due to the powerful multimodal representation capabilities of GPT-4V and GPT-4o, they can directly reason with both visual and textual prompts. However, it is challenging for them to make correct VLN decisions from sequential images. This may be attributed to the inability of VLMs in spatial-aware reasoning. As shown in Table 5, the proposed STMR representation with LLMs significantly outperforms the direct input of RGB images into VLMs. It demonstrates that LLMs can get better spatial reasoning abilities when provided with proper visual encoding.
Grid Size of STMR.
In Fig. 7, we demonstrate the necessity of maintaining a trade-off between grid resolution and the total number of grid cells in STMR representation. Set the resolution to 5m may not allow for as precise spatial inference as 2m, but it is enough for outdoor spatial reasoning.
State Update of Sub-goal.
@@
To ensure that each sub-goal can be executed despite environmental changes, the LLM does not change the original path plan in each iteration. Instead, it updates the status of the current sub-goal, including todo, in process, and completed. We compare this strategy with one that updates the entire multi-step path planning (row 1 of Table 6) based on the observation in each iteration. This strategy does not notably enhance the OSR performance, indicating that the proposed STMR already furnishes enough capacity for navigating the environment comprehensively. Nonetheless, it does positively impact the decision-making process, leading to an increase in the success rate (SR) from 9.0% to 15.0%.
VLM Input RGB STMR NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
@@
$\checkmark$ 88.7 15.0 28.0 3.6
Table 5: Ablation study on the different visual prompts for VLMs.
Figure 7: Ablation study on different STMR grid sizes.
Module NE/m $\downarrow$ SR/% $\uparrow$ OSR/% $\uparrow$ SDTW/% $\uparrow$
@@
w state 88.7 15.0 28.0 3.6
Table 6: Ablation study on the state update strategies.
Conclusions
This paper addresses the challenging aerial VLN task by proposing an LLM-based training-free framework. To enhance the spatial reasoning ability of LLMs, we design Semantic-Topo-Metric Representation (STMR). STMR first integrates instruction-related landmarks and their locations into a top-down map, and subsequently transforms this map into a matrix representation containing semantic, topological, and distance metric information. Taking the proposed STMR as a part of the LLM prompts, we significantly improve the UAV’s navigation capabilities. Our framework achieves state-of-the-art results on both simple navigation tasks and complex navigation tasks, demonstrating its effectiveness and robustness.
Limitations and Future Work
While the proposed STMR significantly improves LLM’s spatial understanding ability, it still struggles to handle very precise distances in 3D space. Additionally, pretrained perception models sometimes introduce errors under different observation angles and distances, causing difficulties for subsequent navigation. Future avenues of improvement include using the primary viewpoint to assist spatial perception, or designing more robust perception strategies.